# SLAC: Simulation-Pretrained Latent Action Space for Whole-Body Real-World RL

### v1

@@ Abstract
Building capable household and industrial robots requires mastering the control of versatile, high-degree-of-freedom (DoF) systems such as mobile manipulators. While reinforcement learning (RL) holds promise for autonomously acquiring robot control policies, scaling it to high-DoF embodiments remains challenging. Direct RL in the real world demands both safe exploration and high sample efficiency, which are difficult to achieve in practice. Sim-to-real RL, on the other hand, is often brittle due to the reality gap. This paper introduces SLAC, a method that renders real-world RL feasible for complex embodiments by leveraging a low-fidelity simulator to pretrain a task-agnostic latent action space. SLAC trains this latent action space via a customized unsupervised skill discovery method designed to promote temporal abstraction, disentanglement, and safety, thereby facilitating efficient downstream learning. Once a latent action space is learned, SLAC uses it as the action interface for a novel off-policy RL algorithm to autonomously learn downstream tasks through real-world interactions. We evaluate SLAC against existing methods on a suite of bimanual mobile manipulation tasks, where it achieves state-of-the-art performance. Notably, SLAC learns contact-rich whole-body tasks in under an hour of real-world interactions, without relying on any demonstrations or hand-crafted behavior priors. More information, code, and videos at robo-rl.github.io
@@ 3 SLAC: Simulation-Pretrained Latent Action Space for Real-World RL
Due to the high dimensionality of and the complexity of real-world tasks, directly optimizing in the real world is prohibitively sample-inefficient and unsafe. To address these issues, we propose to replace the native control space with an -dimensional multi-discrete 222While our method is compatible with both continuous and discrete latent actions, we focus on the discrete case in our experiments. latent action space learned in a low-fidelity simulation, which does not accurately replicate the visual or physical properties of the real world and does not implement the task reward , but approximately retains key physical affordances and shares the same robot action space .
@@ 3.1 Learning a Latent Action Space in Simulation
However, despite its potential, USD has seen limited adoption in robotics due to its high sample complexity, which renders it impractical for direct deployment on real robots. Instead, SLAC addresses this limitation by conducting USD entirely in simulation, where data collection is fast and inexpensive. Our key insight is that even low-fidelity simulation can serve as an effective substrate for behavior pretraining: as long as the learned skills span a sufficiently diverse range of behaviors, they can be composed downstream to solve real-world tasks. To this end, SLAC employs simulation environments that do not directly replicate a real-world counterpart (e.g. not visually realistic, no hard-to-simulate objects like marker traces), but still preserve key geometric affordances that are potentially useful for downstream tasks (e.g. a whiteboard that the robot can touch, an obstacle that the robot may collide with). We show some of these environments in Fig. 1. Importantly, such a simulation environment is often significantly easier to create than a high-fidelity one.
@@
We directly optimize Eq. 5 by running online Reinforcement Learning in simulation. We show the pseudo-code for latent action space learning in the appendix.
@@ 3.2 Sample-Efficient Policy Learning of Downstream Tasks in the Real World
In practice, we implement the Q-decomposition by masking out irrelevant action dimensions for each using a binary adjacency matrix that encodes action-reward dependencies. This implementation enables parallel computation of all Q-functions, significantly accelerating training. The matrix can either be learned automatically from a small number of random interaction trajectories [10], or manually specified when the mapping is known a priori [13]. In SLAC, we assume access to this mapping as a design choice, leveraging the fact that each latent action dimension is intentionally constructed to control a specific environment entity (see Sec. 3.1).
@@ 4 Experimental Results
Our main experiments evaluate SLAC on a bi-manual mobile manipulator – an ideal application domain for our method, which is not only extremely challenging due to the complexity of embodiment and tasks, but also tremendously useful as household robots. Specifically, we evaluate our methods in two different environments: a table environment, where the robot faces a table with objects on it, and a whiteboard environment where the robot can interact with a whiteboard. Each simulation environment takes less than 20 minutes to create in iGibson [61]. In each environment (shown in Fig. 1, described in detail in the appendix), we evaluate multiple different visuomotor contact-rich tasks that require whole-body motion to solve, and present the results (Sec. 4.1) and ablations (Sec. 4.2). Finally, we present additional results in Sec. 4.3 on a multi-agent domain, demonstrating the broad applicability of our method.
@@ 4.1 Training Setup & Results
RLPD [57], a state-of-the-art method for learning from both online and prior data, which we use for training in the real world with prior data from simulation.
@@
Metric and Results: For each method on each downstream task, we compare the success rate of the final policy across 10 different trials. For the three methods that require training in simulation (SLAC, Sim2Real, RLPD), we train for 10M steps in simulation for each task. For the three methods that require real-world interactions (SLAC, SERL, RLPD), we train each of them for 30k steps of real-world low-level robot actions, corresponding to 50 minutes of real-world interactions, and additionally report the number of times they have violated the safety constraints during training. The full result is shown in Table 1. In all four tasks, SLAC learns to solve the task in less than an hour of real-world interactions (curves in Fig. 3(a)), while maintaining safety during real-world exploration, significantly outperforming the baselines.
@@ 4.2 Ablations
Not Temporally Extended: where the task policy makes decisions at the same frequency as the latent action encoder (i.e. 10hz).
@@ 5 Conclusion
Limitations: Despite its strong empirical performance, SLAC is not without limitations. First, SLAC introduces an implicit trade-off related to the granularity of the latent action space. For example, an identity mapping between the robot’s raw action space and latent action space would make the downstream task policy capable of learning any task within the original capability of the robot, but would significantly reduce the sample efficiency (as shown in our results in Sec. 4.1). Similarly, the temporal length of each latent action entails another tradeoff, where shorter latent actions give more control to the task policy at the expense of longer task horizons, which may hamper learning (as shown in our ablation studies in Sec. 4.2). It is likely that the optimality of the latent action space will be strongly task-dependent. Second, in the current SLAC framework, the latent action decoder is kept fixed during downstream learning. However, for more fine-grained tasks, we might benefit from finetuning the latent action encoder while training the task policy. While this is conceptually doable via the option framework [64], we leave the empirical study of how to implement it in a stable manner to future work. Finally, while this paper primarily focuses on the algorithmic side of real-world learning, we expect future engineering efforts in the automation of task reset and downstream reward generation (e.g., via VLM/LLM or a learned reward function) to further boost the downstream learning efficiency.
@@ A.1 Policy and Training Visualizations
A.2 Q Decomposition Proof
@@
A.3 Universal Safety Reward
@@
A.4 SLAC Unsupervised Latent Action Space Learning Pseudocode
@@
A.5 SLAC Real-World Downstream RL Pseudocode
@@
A.6 Hyperparameter
Here, we present the hyperparameter for both the latent action decoder training and the downstream task learning. The same hyperparamters are shared across all tasks.
@@ A.6 Hyperparameter
A.7 Environment Description
@@ A.7.2 Table Environment
A.8 Discussion of Broader Areas Related to SLAC

### v3

@@ Abstract
Building capable household and industrial robots requires mastering the control of versatile, high-degree-of-freedom (DoF) systems such as mobile manipulators. While reinforcement learning (RL) holds promise for autonomously acquiring robot control policies, scaling it to high-DoF embodiments remains challenging. Direct RL in the real world demands both safe exploration and high sample efficiency, which are difficult to achieve in practice. Sim-to-real RL, on the other hand, is often brittle due to the reality gap. This paper introduces SLAC, a method that renders real-world RL feasible for complex embodiments by leveraging a low-fidelity simulator to pretrain a task-agnostic latent action space. SLAC trains this latent action space via a customized unsupervised skill discovery method designed to promote temporal abstraction, disentanglement, and safety, thereby facilitating efficient downstream learning. Once a latent action space is learned, SLAC uses it as the action interface for a novel off-policy RL algorithm to autonomously learn downstream tasks through real-world interactions. We evaluate SLAC against existing methods on a suite of bimanual mobile manipulation tasks, where it achieves state-of-the-art performance. Notably, SLAC learns contact-rich whole-body tasks in under an hour of real-world interactions, without relying on any demonstrations or hand-crafted behavior priors. More information and robot videos at robo-rl.github.io
@@ 3 SLAC: Simulation-Pretrained Latent Action Space for Real-World RL
Due to the high dimensionality of and the complexity of real-world tasks, directly optimizing in the real world is prohibitively sample-inefficient and unsafe. To address these issues, we propose to replace the native control space with an -dimensional multi-discrete 222While our method is compatible with both continuous and discrete latent actions, we focus on the discrete case in our experiments. We discuss this choice in the appendix. latent action space learned in a low-fidelity simulation, which does not accurately replicate the visual or physical properties of the real world and does not implement the task reward , but approximately retains key physical affordances and shares the same robot action space .
@@ 3.1 Learning a Latent Action Space in Simulation
However, despite its potential, USD has seen limited adoption in robotics due to its high sample complexity, which renders it impractical for direct deployment on real robots. Instead, SLAC addresses this limitation by conducting USD entirely in simulation, where data collection is fast and inexpensive. Our key insight is that even low-fidelity simulation can serve as an effective substrate for behavior pretraining: as long as the learned skills span a sufficiently diverse range of behaviors, they can be composed downstream to solve real-world tasks. To this end, SLAC employs simulation environments that do not directly replicate a real-world counterpart (e.g. not visually realistic, no hard-to-simulate objects like marker traces), but still preserve key geometric affordances that are potentially useful for downstream tasks (e.g. a whiteboard that the robot can touch, an obstacle that the robot may collide with). We show some of these environments in Fig. 1. Importantly, because we do not need the simulation environment to exactly match the real world, prototyping a new environment and training a corresponding latent action space can be done quickly.
@@
We directly optimize Eq. 5 by running online RL in simulation. In appendix, we show the pseudo-code for latent action space learning, and further discuss the properties of the learned action space.
@@ 3.2 Sample-Efficient Policy Learning of Downstream Tasks in the Real World
In practice, we implement the Q-decomposition by masking out irrelevant action dimensions for each using a binary adjacency matrix that encodes action-reward dependencies. This implementation enables parallel computation of all Q-functions, significantly accelerating training. The matrix can either be learned automatically from a small number of random interaction trajectories [10], or manually specified when the mapping is known a priori [13]. In SLAC, we assume ground-truth access to this mapping, leveraging the fact that each latent action dimension is intentionally constructed to control a specific environment entity (see Sec. 3.1).
@@ 4 Experimental Results
Our primary experiments evaluate SLAC on a bi-manual mobile manipulator – a domain that is ideally suited to our approach. This domain is not only highly challenging due to the complexity of the embodiment and task space, but also practically important for the development of capable household robots. Specifically, we evaluate our methods in two different environments: a table environment, where the robot faces a table with objects on it, and a whiteboard environment where the robot can interact with a whiteboard. Each simulation environment takes less than 20 minutes to create in iGibson [61], using off-the-shelf object models without any real2sim. In each environment (shown in Fig. 1, described in detail in the appendix), we evaluate multiple different visuomotor contact-rich tasks that require whole-body motion to solve, and present the results (Sec. 4.1) and ablations (Sec. 4.2). Finally, we present additional results in Sec. 4.3 on a multi-agent domain, demonstrating the broad applicability of our method.
@@ 4.1 Training Setup & Results
RLPD [57], a state-of-the-art method for learning from both online and prior data, which we use for finetuning in the real world with prior data from simulation.
@@
Metric and Results: For each method on each downstream task, we compare the success rate of the final policy across 10 rollouts with different initial states. For the three methods that require training in simulation (SLAC, Sim2Real, RLPD), we train for 10M steps in simulation for each task. For the three methods that require real-world interactions (SLAC, SERL, RLPD), we train each of them for 30k steps of real-world low-level robot actions, corresponding to 50 minutes of real-world interactions, and additionally report the number of times they have violated the safety constraints during training. The full result is shown in Table 1. In all four tasks, SLAC learns to solve the task in less than an hour of real-world interactions (curves in Fig. 3(a)), while maintaining safety during real-world exploration, significantly outperforming the baselines.
@@ 4.2 Ablations
Not Temporally Extended: where the task policy makes decisions at the same frequency as the latent action decoder (i.e. 10hz).
@@ 5 Conclusion
Limitations: Despite its strong empirical performance, SLAC is not without limitations. First, SLAC introduces an implicit trade-off related to the granularity of the latent action space. For example, an identity mapping between the robot’s raw action space and latent action space would make the downstream task policy capable of learning any task within the original capability of the robot, but would significantly reduce the sample efficiency (as shown in our results in Sec. 4.1). Similarly, the temporal length of each latent action entails another tradeoff, where shorter latent actions give more control to the task policy at the expense of longer task horizons, which may hamper learning (as shown in our ablation studies in Sec. 4.2). It is likely that the optimality of the latent action space will be strongly task-dependent. Second, in the current SLAC framework, the latent action decoder is kept fixed during downstream learning. However, for more fine-grained tasks, we might benefit from finetuning the latent action decoder while training the task policy. While this is conceptually doable via the option framework [64], we leave the empirical study of how to implement it in a stable manner to future work. Finally, while this paper primarily focuses on the algorithmic side of real-world learning, we expect future engineering efforts in the automation of task reset and downstream reward generation (e.g., via VLM/LLM or a learned reward function) to further boost the downstream learning efficiency.
@@ A.1 Policy and Training Visualizations
A.2 Properties of the Learned Latent Action Space
Here, we discuss in detail the properties of the learned SLAC latent action space. In short, the latent action space of SLAC is environment-aware but task-agnostic. It is task-agnostic because it is trained without a task reward, and is only encouraged to induce diversity in behavior following the USD objective. Therefore, the same latent action space can tackle different tasks within a particular environment (e.g., the “push to tray” and “sweep to bag” tasks in our experiment utilize the same latent action space). On the other hand, the latent action space is environment-aware because it is trained to induce diverse behavior in a particular scene in simulation. Note that our latent action space is robust to small variance in the environment: for example, the action space learned in the board environment can be used to learn policies that wipe marks and avoid the trash can, even though there is neither a trash can nor wipeable marks in the simulated board environment.
A.3 Q Decomposition Proof
@@
A.4 Universal Safety Reward
@@
A.5 SLAC Unsupervised Latent Action Space Learning Pseudocode
@@
A.6 SLAC Real-World Downstream RL Pseudocode
@@
A.7 Hyperparameter
Here, we present the hyperparameter for both the latent action decoder training and the downstream task learning. The same hyperparamters are shared across all tasks. We use a low-level step size for all our experiments.
@@ A.7 Hyperparameter
A.8 Discrete vs Continuous Latent Actions
The SLAC framework supports both discrete and continuous latent action. In our experiments, we made a deliberate choice to use a discrete latent action space since a discrete latent action space encodes a compact set of distinguishable behaviors, making it more amenable to hierarchical downstream RL. This is also a standard design choice of previous unsupervised skill discovery methods [22, 25, 24].
A.9 Environment Description
@@ A.9.2 Table Environment
A.10 Discussion of Broader Areas Related to SLAC